# Symmetrical Visual Contrastive Optimization: Aligning Vision-Language Models with Minimal Contrastive Images

### v1

@@ 2 Preliminaries
Our visual contrastive objective is inspired by Direct Preference Optimization (DPO) (Rafailov et al., 2024) that contrasts pairs of textual responses. When applied to VLMs, DPO incorporates the image as an additional prefix condition (Zhou et al., 2024; Li et al., 2023a; Yu et al., 2024; Sarkar et al., 2024). Let be the policy VLM to be optimized, and the fixed reference model (typically the unfinetuend VLM in DPO’s framework) used to measure and optimize how the finetuned policy improves. Given a query (an instruction prompt), an image , and a pair of responses – one preferred (winning) and one dispreferred (losing) – the DPO objective for VLMs can be formulated as:
@@
The above visual conditional preference formulation shifts the target of preference to the images, where the original image is always preferred over the corrupted version . However, this approach shares similar limitations with DPO, as it prioritizes distinguishing image differences without necessarily grounding those differences in the associated texts. Since is typically a noisy variant of with no definitive relationship to , the model could easily rely on superficial visual features to discern the paired images. This encourages shortcut learning, where the model rejects “unrealistic” images like without examining the visual details related to the text tokens.
@@ 4.1 Visual Counterfactual Data
To complement S-VCO, we introduce MVC: a dataset of paired visual contrastive samples designed to enhance the model’s ability to discern visual details. Built on existing sources (e.g., Zhang et al., 2024; Liu et al., 2024c; Gaur et al., 2024), MVC contains image pairs with minimal but meaningful variations, accompanied by corresponding contrastive texts. The curated data includes four key contrast types, as shown in part a). of Figure 4: Object Replacement (changing a specific object); Attribute Replacement (modifying an object’s features like color, shape, or size); Count Modification (altering the number of objects); Position Flipping (reversing relative positions of objects). Except for the last type, these counterfactual images are generated through controlled image synthesis, including inpainting, editing, and generation (Li et al., 2023b; Zhuang et al., 2025; Betker et al., ). In this work, we use CounterCurate (Zhang et al., 2024) and FineCops-Ref (Liu et al., 2024c) as our visual counterfactual data sources. Refer to Appendix A for more dataset statistics.
@@ 4.2 Filtering and Language Augmentation
Language Augmentation. Although visual counterfactual data sources provide image contrasts, their original textual descriptions are typically short captions, which are unideal for VLM finetuning. To address this, we augment the data using a two-step process with a strong LLM (gpt-4o), as shown in part c). of Figure 4: 1. Generating queries: For each pair of captions, we prompt the LLM to generate a conversational question as if the captions were natural responses to that question, while ensuring the contrasts remain explicit. 2. Rewriting and diversifying: We prompt the LLM to rephrase the original captions lacking clarity on key contrasts, thus enabling a more natural flow of language given the generated question and the contrastive details. Overall, the augmentation step results in conversational instruction-response pairs that are more suited for VLM finetuning, and more aligned with the visual details in the contrasting image pairs. Our prompt template is provided in Appendix C.
@@ 5.1 Setup
Base VLMs. We use two pretrained models hosted on Huggingface: LLaVA-1.5-7B222llava-hf/llava-1.5-7b-hf (Liu et al., 2024a) – denoted as LV-1.5(-7B), and LLaVA-Next-Interleave-7B333llava-hf/llava-interleave-qwen-7b-hf (Li et al., 2024) – denoted as LV-INT(-7B).
@@ 5.2 Main Results
S-VCO achieves consistently superior performance across benchmarks in various domains. Figure 1 illustrates performance improvements over the base-VLM (LV-INT) for each benchmark category. Our S-VCO consistently outperforms baseline methods across nearly all domains, with only a slight drop in knowledge-heavy tasks like ScienceQA that relies minimally on visual input (discussed more below). The most significant gain is in visual hallucination tasks, where S-VCO enhances the base model by over %. Considerable improvements are also observed in vision-centric (%) and general domains (%). Compared to baseline methods, DPO and mDPO trained on the standard instruction-tuning dataset VLF, S-VCO with MVC delivers much greater and more consistent improvements across domains. While recent approaches like mDPO (Wang et al., 2024) improve visual hallucination metrics, its effects on other abilities are less notable. In contrast, S-VCO not only excels in visually demanding tasks but also achieves considerable gains in other domains.
S-VCO shows increasing benefits as benchmarks become more visually dependent. We quantify visual dependency of a metric as the percentage drop in a base-VLM’s performance when image inputs are removed. Using LV-INT as the base model, we rank benchmarks by visual dependency and plot the improvement trends of different methods in Figure 6. Dotted lines show fitted regressions, and shaded areas represent variances. S-VCO demonstrates increasingly pronounced improvements as visual dependency rises, aligning with its design focus on strengthening visual detail recognition (Section 3). For visually highly dependent benchmarks like LLaVABench (Liu et al., 2024b), MMVet (Yu et al., 2023), and MM-Hal (Sun et al., 2023), S-VCO delivers the most substantial gains. In contrast, ScienceQA (SQA) (Lu et al., 2022), which benefits merely ~% from image inputs, sees a minor drop in performance, as visual information plays little role in this task. Compared to other methods, SFT (Section 5.3) degrades performance on visually dependent metrics, while other preference tuning methods show positive trends. Among all, S-VCO achieves the most significant trend of gains with increasingly vision-intensive tasks.
@@ 5.3 Ablations
Comparing to SFT. To investigate whether standard supervised finetuning (SFT) on both sides of the image-text pairs could achieve similar results, we construct a new instruction-tuning dataset by including both and from MVC, effectively doubling its size (). Results for SFT are shown in the last row of Table 1. While SFT preserves performance on ScienceQA – a benchmark minimally reliant on visual inputs – it performs significantly worse across all other domains, achieving the lowest results among all methods. Unlike S-VCO, SFT lacks the contrastive supervision necessary to highlight subtle visual-textual alignments, thus leading to poor performance on tasks requiring strong visual grounding.
@@ 6 Related Work
VLM’s Visual Hallucinations. Recent studies (Dai et al., 2024; Deng et al., 2024; Tong et al., 2024b; Chen et al., 2024; Jia et al., 2024) observed the tendency of VLMs to hallucinate content that is not present in the visual input, indicating a lack of robust multimodal grounding. To address this issue, several training-free methods have been proposed. Yang et al. (2023) introduced the “Set-of-Mark” prompting technique, which overlays spatial and textual markers on images, helping models like GPT-4V better reference specific regions. Deng et al. (2024) employed CLIP-guided decoding to steer the language outputs with grounded visual cues. On the architectural side, GRILL Jin et al. (2023) incorporates object-level alignment during pretraining to prompt visual grounding. In contrast to these approaches, our work focuses on finetuning with a novel preference-tuning objective (Section 3) and a data construction pipeline (Section 4) based on visual counterfactuals (Zhang et al., 2024; Liu et al., 2024c), targeting more precise alignment between multimodal details.
VLM Finetuning. Finetuning enhances VLMs for task-specific performance and alignment with human preferences. SFT remains widely adopted to guide models toward instruction-following behaviors in multimodal contexts (Sun et al., 2024; Jiang et al., 2024b). DPO (Rafailov et al., 2024) optimizes the margin between the finetuned and unfintuned model versions using paired preference data. Extensions of DPO to VLMs incorporate the image as an additional prefix condition (Zhou et al., 2024). Recent methods such as mDPO (Wang et al., 2024), MFPO (Jiang et al., 2024a), V-DPO (Xie et al., 2024), CHiP (Fu et al., 2025) and Image-DPO (Luo et al., 2024) further adapt the preference tuning paradigm to focus on image-side preferences over a pair of “good” and “bad” image, aiming to reduce visual hallucinations. Our approach S-VCO discards the one-sided “preference” formulation by introducing a stricter visual contrastive objective within a symmetrical construct, in which way the “preference” is treated as alignment over a matching image-text pair. This enables more comprehensive and robust improvements in VLM performance across tasks.

### v2

@@ 2 Preliminaries
Our visual contrastive objective is inspired by Direct Preference Optimization (DPO) (Rafailov et al., 2024) that contrasts pairs of textual responses. When applied to VLMs, DPO incorporates the image as an additional prefix condition (Zhou et al., 2024; Li et al., 2023a; Yu et al., 2024; Sarkar et al., 2024). Let be the policy VLM to be optimized, and the fixed reference model (typically the unfinetuend VLM in DPO’s framework) used to measure how the finetuned policy improves. Given a query (an instruction prompt), an image , and a pair of responses – one preferred (winning) and one dispreferred (losing) – the DPO objective for VLMs can be formulated as:
@@
The above visual conditional preference formulation shifts the target of preference onto the images, where the original image is always preferred over the corrupted version . However, this approach shares similar limitations with DPO, as it prioritizes distinguishing image differences without necessarily grounding those differences in the associated texts. Since is typically a noisy variant of with no definitive relations to , the model could easily rely on superficial visual features to discern the image pair. This encourages shortcut learning, where the model rejects “unrealistic” images like without examining the visual details related to the text tokens.
@@ 4.1 Visual Counterfactual Data
To complement S-VCO, we introduce MVC: a dataset of paired visual contrastive samples designed to enhance the model’s ability to discern visual details. Built on existing sources (e.g., Zhang et al., 2024; Liu et al., 2024c; Gaur et al., 2024), MVC contains image pairs with minimal but meaningful variations, accompanied by corresponding contrastive texts. The curated data includes four key contrast types, as shown in part a). of Figure 4: Object Replacement (changing a specific object); Attribute Replacement (modifying an object’s features like color, shape, or size); Count Modification (altering the number of objects); Position Flipping (reversing relative positions of objects). Except for the last type, these counterfactual images are generated through controlled image synthesis, including inpainting, editing, and generation (Li et al., 2023b; Zhuang et al., 2025; Betker et al., ). In this work, we use CounterCurate (Zhang et al., 2024) and FineCops-Ref (Liu et al., 2024c) as our visual counterfactual data sources. Refer to Appendix A for dataset statistics.
@@ 4.2 Filtering and Language Augmentation
Language Augmentation. Although visual counterfactual data sources provide image contrasts, their original textual descriptions are typically short captions, which are unideal for VLM finetuning. To address this, we augment the data using a two-step process with a strong LLM (gpt-4o), as shown in part c). of Figure 4: 1. Generating queries: For each pair of captions, we prompt the LLM to generate a conversational question as if the captions were natural responses to that question, while ensuring that the contrasts remain explicit. 2. Rewriting and diversifying: We prompt the LLM to rephrase the original captions lacking clarity on key contrasts, thus enabling a more natural flow of language given the generated question and the contrastive details. Overall, the augmentation step results in conversational instruction-response pairs that are more suited for VLM finetuning, and more aligned with the visual details in the contrasting image pairs. Our prompt templates are provided in Appendix C.
@@ 5.1 Setup
Base VLMs. We use two pretrained models hosted on Huggingface: LLaVA-1.5-7B (Liu et al., 2024a) – denoted as LV-1.5(-7B), and LLaVA-Next-Interleave-7B (Li et al., 2024b) – denoted as LV-INT(-7B).
@@ 5.2 Main Results
S-VCO achieves consistently superior performance across benchmarks in various domains. Figure 1 illustrates performance improvements over the base-VLM (LV-INT) for each benchmark category. Our S-VCO consistently outperforms baseline methods across nearly all domains, with only a slight drop in knowledge-heavy tasks like ScienceQA that relies minimally on visual input (discussed more below). The most significant gain is in visual hallucination tasks, where S-VCO enhances the base model by ~%. Considerable improvements are also observed in vision-centric (~%) and general domains (~%). Compared to baseline methods, DPO and mDPO trained on the standard instruction-tuning dataset VLF, S-VCO with MVC delivers much greater and more consistent improvements across domains. While recent approaches like mDPO (Wang et al., 2024) improve visual hallucination metrics, its effects on other abilities are less notable. In contrast, S-VCO not only excels in visually demanding tasks but also achieves considerable gains in other domains.
S-VCO shows increasing benefits as benchmarks become more visually dependent. We quantify visual dependency of a metric as the percentage drop in a base-VLM’s performance when image inputs are removed. Using LV-INT as the base model, we rank benchmarks by visual dependency (e.g., MMVP accuracy drops to without image inputs), and plot the improvement trends of different methods in Figure 6. Dotted lines show fitted regressions, and shaded areas represent variances. S-VCO demonstrates increasingly pronounced improvements as visual dependency rises, aligning with its design focus on strengthening visual detail recognition (Section 3). On highly visually-dependent benchmarks such as MM-Hal, LLaVABench, MMVet and MMVP, our S-VCO delivers the most substantial gains. In contrast, on SienceQA (SQA), which benefits merely ~% from image inputs, we observe a minor drop in performance after S-VCO, as visual information plays very little role in this task. Compared to other methods, SFT degrades performance on visually dependent metrics (Section 5.3), while preference tuning approaches show positive trends. Among all, S-VCO achieves the most significant trend of gains with increasingly vision-intensive tasks.
@@ 5.3 Ablations
Comparing to SFT. To investigate whether standard supervised finetuning (SFT) on both sides of the image-text pairs could achieve similar results, we construct a new instruction-tuning dataset by including both and from MVC, effectively doubling its size (). Results for SFT are shown in the last row of Table 1. While SFT preserves performance on ScienceQA – a benchmark minimally reliant on visual inputs – it performs significantly worse across all other domains, yielding the lowest results among all methods. Unlike S-VCO, SFT lacks the contrastive supervision necessary to highlight subtle visual-text alignment, thus leading to poor performance on tasks requiring strong visual grounding.
@@ 6 Related Work
VLM’s Visual Hallucinations. Recent studies (Deng et al., 2024; Tong et al., 2024b; Chen et al., 2024; Jia et al., 2024) have shown that VLMs tend to hallucinate content not present in the visual input. VLMs also struggle with fine-grained visual understanding (e.g., recognizing object attributes and relations) – especially when tested on confounding image pairs (Peng et al., 2024; Li et al., 2024a). This aligns with our findings in Figure 2, suggesting a lack of robust and faithful multimodal grounding. To address this issue, several training-free methods have been proposed. Yang et al. (2023) introduced “Set-of-Mark” prompting that overlays spatial and textual markers on images to help models reference specific regions. Deng et al. (2024) employed CLIP-guided decoding to steer the language outputs with grounded visual cues. Architecture-wise, GRILL Jin et al. (2023) incorporates object-level alignment during pretraining to promote visual grounding. Unlike previous approaches, our work focuses on finetuning with a novel objective (Section 3) and a data construction pipeline (Section 4) based on visual counterfactuals (Zhang et al., 2024; Liu et al., 2024c), targeting more precise alignment of multimodal details.
VLM Finetuning. Finetuning improves task-specific performance of VLMs and aligns them better with human preferences. SFT remains widely adopted to guide models toward towards following instructions (Sun et al., 2024; Jiang et al., 2024b). DPO (Rafailov et al., 2024) optimizes the margin between finetuned and unfinetuned model versions using paired preference data. Its extension to VLMs incorporates image as additional prefix condition (Zhou et al., 2024). Recent methods such as mDPO (Wang et al., 2024), MFPO (Jiang et al., 2024a), V-DPO (Xie et al., 2024), CHiP (Fu et al., 2025) and Image-DPO (Luo et al., 2024) further adapt the preference tuning paradigm to focus on image-side preferences over a pair of “good” and “bad” image, aiming to reduce visual hallucinations. Our approach S-VCO replaces the one-sided “preference” formulation with a stricter visual contrastive objective of symmetrical construct, treating “preference” explicitly as alignment over matching image-text pairs. This enables more comprehensive and robust VLM improvements across tasks.